### v2

# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

Abstract
Since multiple MRI contrasts of the same anatomy contain redundant information, one contrast can be used as a prior for guiding the reconstruction of an undersampled subsequent contrast. To this end, several learning-based guided reconstruction methods have been proposed. However, a key challenge is the requirement of large paired training datasets comprising raw data and aligned reference images. We propose a modular two-stage approach for guided reconstruction addressing this issue, which additionally provides an explanatory framework for the multi-contrast problem in terms of the shared and non-shared generative factors underlying two given contrasts. A content/style model of two-contrast image data is learned from a largely unpaired image-domain dataset and is subsequently applied as a plug-and-play operator in iterative reconstruction. The disentanglement of content and style allows explicit representation of contrast-independent and contrast-specific factors. Based on this, incorporating prior information into the reconstruction reduces to simply replacing the aliased content of the image estimate with high-quality content derived from the reference scan. Combining this component with a data consistency step and introducing a general corrective process for the content yields an iterative scheme. We name this novel approach PnP-MUNIT. Various aspects like interpretability and convergence are explored via simulations. Furthermore, its practicality is demonstrated on the NYU fastMRI DICOM dataset and two in-house multi-coil raw datasets, obtaining up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM. In a radiological task, PnP-MUNIT allowed 33.3% more acceleration over clinical reconstruction at diagnostic quality.
I Introduction
Magnetic resonance imaging (MRI) is an invaluable medical imaging modality due to the high-quality scans it delivers, the variety of complementary information it can capture, and its lack of radiation-related risks, leading to its wide usage in clinical practice. However, its central limitation is the inherently slow data acquisition process. The raw sensor data is acquired in the frequency domain, or k-space, from which the image is reconstructed. Over the last 25 years, advancements such as parallel imaging [1, 2], compressed sensing (CS) [3], and deep learning reconstruction [4, 5] have enabled considerable speedups by allowing sub-Nyquist k-space sampling and relying on computationally sophisticated reconstruction. These techniques have subsequently been implemented on commercial MRI systems and have demonstrated (potential) improvements in clinical workflow [6].
A clinical MRI session typically involves the acquisition of multiple scans of the same anatomy through the application of different MR pulse sequences, exhibiting different contrasts. Because these scans are different reflections of the same underlying reality, they share a high degree of shared structure. However, currently deployed clinical protocols acquire and reconstruct each scan as an independent measurement, not leveraging the information redundancy across scans. There is, therefore, an opportunity to further optimize MRI sessions by exploiting this shared information. On the reconstruction side, multi-contrast methods have addressed this problem by introducing the shared information into the reconstruction phase to allow higher levels of k-space undersampling. In the simplest case of two contrasts, multi-contrast reconstruction can be classified into two types – (a) guided reconstruction, where an existing high-quality reference scan is used to guide the reconstruction of an undersampled second scan [7, 8, 9] and (b) joint reconstruction, where both contrasts are undersampled and are reconstructed simultaneously [10, 11, 12]. In this work, we consider the problem of guided reconstruction, assuming no inter-scan motion between reference and target scans.
Figure 1: Our two-stage approach to guided reconstruction. (a) The first stage learns a content/style model of two-contrast MR image data. The two image domains $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ are decomposed into a shared content domain $\mathcal{C}$ and separate style domains $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. This illustration only provides an intuition about the information encoded in the content and style domains, and does not represent the actual training process, which is given in detail in Section III-B. (b) The reconstruction stage applies this model as a content consistency operator (bottom) within an ISTA-based iterative algorithm. Given an aligned reference image, guidance is introduced into the reconstruction by simply replacing its aliased content with content derived from the reference. CR denotes a content refinement update, which iteratively corrects for inconsistencies between the reference content and the measured k-space data, improving the effectiveness of the content consistency operator. DC denotes data consistency. Wavelet-domain soft-thresholding (top) and CNN-based denoising (middle) used in CS-WT and PnP-CNN reconstruction, respectively, are shown for comparison.
The guided reconstruction problem entails using the local structure of the reference scan as a prior to complement the undersampled k-space measurements of the target scan. This problem has been formulated in different ways, ranging from conventional CS [7, 8] to end-to-end learning with unrolled networks [9, 13, 14, 15] and, more recently, diffusion model-based Bayesian maximum a posteriori estimation [16]. Most end-to-end approaches, although more powerful than earlier hand-crafted ones, suffer from the main drawback of requiring large paired training datasets consisting of the target image and its k-space together with an aligned reference image, thereby limiting their application on real-world MR data. We address this issue by proposing a plug-and-play reconstruction method that splits the problem into a purely image-domain learning sub-problem and an iterative reconstruction sub-problem. The learning problem leverages ideas from content/style decomposition, thereby offering a degree of interpretability to our approach. To the best of our knowledge, ours is the first work that thoroughly explores content/style-based generative modeling for multi-contrast MRI reconstruction.
In recent years, image-to-image translation has found application in the direct estimation of one MR contrast from another [17, 18, 19, 20]. While these methods are attractive due to their lack of dependence on k-space data, viewed in the light of MR physics, cross-contrast translation takes an extreme stance by not explicitly taking into account contrast-specific sensitivities and relying solely on the prior contrast. Hence, in the context of MR image formation, cross-contrast translation, by itself, can only provide a part of the information about the target image as data acquisition via MR sequences is necessary to obtain new information about the anatomy. That being said, literature on unpaired image translation provides a repertoire of useful tools such as joint generative modeling of two-domain image data [21, 22], which can be adopted to complement image reconstruction. We observe that methods such as MUNIT [22] can be applied to learn semantically meaningful representations of contrast-independent and contrast-specific information as content and style, respectively, without the need for paired image-domain training data.
Plug-and-play (PnP) methods are an emerging paradigm for solving inverse problems in computational imaging. The main research line [23, 24] has focused on learning CNN-based denoising models on image-domain data and applying them as functions replacing proximal operators in iterative algorithms like ISTA and ADMM, demonstrating improved image recovery. An advantage of this approach is the decoupling of the learning problem of image modeling from the inverse problem of image reconstruction, thereby simplifying model training and improving generalizability across different acceleration factors, undersampling patterns, etc. With this design pattern in mind, we combine content/style image modeling and iterative reconstruction in a PnP-like framework.
We first leverage semantic content/style modeling to learn explicit representations of contrast-independent and contrast-specific components from two-contrast MR image data. This training process is independent of the reconstruction problem and can be performed using unpaired images. We then make the interesting observation that in multi-contrast MR images, the style information tends to localize in the center of the k-space, allowing for an accurate style estimation from an undersampled image. Since the content of the reconstructing image, which is the remaining piece of information and is contrast-independent, is supplied by the reference image, one can compose a de-aliased estimate of the reconstruction from the undersampled image in a single step. We term this the content consistency operation, which forms the basis of our iterative reconstruction algorithm PnP-MUNIT [25]. An overview of our approach is shown in Fig. 1. While the PnP-based decoupling of image modeling and image reconstruction simplifies the training process and allows the two stages to be analyzed separately, a further level of decoupling offered by content/style disentanglement offers additional modularity and an intuitive guidance mechanism for the reconstruction.
Specifically, our contributions are four-fold:
  1. 1.
    We show that unpaired image-domain training can be used to learn disentangled contrast-independent and contrast-specific representations, followed by a fine-tuning strategy that refines the content representation using a modest amount of paired image-domain data.
  2. 2.
    With this content/style model as the basis, we define a content consistency operator capable of removing severe undersampling artifacts from the reconstructing image, given the corresponding reference image.
  3. 3.
    Developing this idea further and incorporating a corrective process for the content, we propose PnP-MUNIT, a modular algorithm for guided reconstruction, combining the flexibility of the plug-and-play approach with the semantic interpretability of content/style decomposition.
  4. 4.
    Through comprehensive experiments, we shed light on several properties of PnP-MUNIT such as convergence and robustness and demonstrate its applicability on real-world raw data and its potential clinical utility for a specific radiological task.
II Related Work
II-A Reconstruction Methods for Accelerated MRI
Techniques for accelerating MRI by k-space undersampling go back to compressed sensing (CS) [3], which combines random sampling with sparsity-based iterative denoising, most commonly implemented based on the ISTA [26] or ADMM [27] family of algorithms. Most modern deep learning-based reconstruction methods focus on improving the denoising part. Plug-and-play (PnP) methods [23, 24] replace the proximal operator in ISTA and ADMM with off-the-shelf denoisers such as a learned convolutional denoising model. Unrolled networks [28, 4, 5] extend this idea by casting the entire iterative algorithm into one large network, trained end-to-end. This makes them more adaptive to factors such as sampling pattern and acceleration, although at a cost of generalizability [23].
One of the earliest CS-based guided reconstruction methods was proposed by Ehrhardt and Betcke [7] introducing structure-guided total variation (STV), which assumes the sparse coefficients of the reconstruction to be partially known based on the edge features in the reference scan. Later work introduced adaptive elements into the multi-contrast CS framework, e.g. Weizman et al. [8] proposed adaptive weighting-based guided CS and Song et al. [29] used adaptive sparse domains based on coupled dictionary learning. In the latter, the problem was formulated using a patch-level linear model comprising coupled and distinct sparse dictionary representations of the two contrasts. This model resembles a content/style model in form, although it is more restrictive. A general drawback of classical methods compared to deep learning-based ones is their lower flexibility. End-to-end learning-based methods [9, 13, 30, 31, 32, 15], on the other hand, supply the reference scan as an additional input to a deep reconstruction model allowing it to automatically learn the suitable features to extract and transfer into the reconstruction, and have proven to be more effective than conventional algorithms. However, end-to-end methods require large paired training datasets comprising the ground-truth image and the k-space together with aligned reference images, which is too strong a constraint when working with retrospectively collected clinical data that reflects the natural inconsistencies of routine practice such as missing contrasts, differences in spatial resolutions, and the existence of inter-scan motion. Additionally, by relying on end-to-end-learned features, these methods are generally less interpretable than their conventional counterparts in that they do not explicitly model the multi-contrast problem in terms of the underlying shared and non-shared information. MC-VarNet [15] is an important exception, which uses a simple linear decomposition of the reference contrast into common and unique components, applying the common component for guidance.
A reconstruction method that is as effective as the learning-based methods while having more lenient data requirements and offering a high degree of interpretability is still needed. Our plug-and-play method relies on learning a deep non-linear content/style transform from image-domain data only, even if subject-wise paired images are not fully available.
II-B Unpaired Image-to-Image Modeling
Image-to-image modeling is the general problem of learning a mapping between two image domains and was first addressed by Pix2Pix [33] and CycleGAN [34] in paired and unpaired settings, respectively. Another line of unpaired image translation methods, the first of which was UNIT [21], assumes a shared latent space underlying the two domains to explicitly represent shared information. However, both CycleGAN and UNIT assume a deterministic one-to-one mapping between the two domains, ignoring the fact that an image in one domain can have multiple valid renderings in the other. Deterministic image translation has been widely applied to MRI contrast-to-contrast synthesis [17, 35, 18]. However, fundamentally, these methods do not account for the variability of the scanning setup that influences the realized contrast level and the differential visibility of pathologies in the target image. Denck et al. [19] partially address this problem by proposing contrast-aware MR image translation where the acquisition sequence parameters are fed into the network to control the output’s contrast level. However, this model is too restrictive since it assumes a single pre-defined mode of variability (i.e. global contrast level) in the data, for which the labels (i.e. sequence parameters) must be available.
MUNIT [22] extended UNIT by modeling domain-specific variability in addition to domain-independent structure, enabling many-to-many mapping and thus overcoming the rigidity of UNIT. The result was a stochastic image translation model which, given an input image, generates a distribution of synthetic images sharing the same “content” but differing in “style”. More fundamentally, MUNIT is a learned invertible transformation between the image domains and the disentangled content/style domains. And unlike other content/style modeling frameworks such as [36, 37, 38], MUNIT models content and style as latent generative factors of the two image domains, providing precise distribution-level definitions for them. Clinical MR images of a given protocol contain multiple modes of variability, many of which are not known a priori. Compared to [19], we make a broader assumption that the contrast-independent semantic information is local in nature and that the contrast-specific variations in the dataset can include global effects of acquisition settings as well as local anatomical features unique to the contrast. We model these using MUNIT, referring to the shared and non-shared components as content and style, respectively.
II-C Combining Image Translation with Reconstruction
Acknowledging the limitation of MR cross-contrast prediction, some prior work has attempted combining it with multi-contrast reconstruction. A naive form of joint image synthesis and reconstruction, e.g. PROSIT [39] and rsGAN [40], involves generating a synthetic image from the reference scan via deterministic image translation and using it in a classical L2-regularized least-squares reconstruction. More recent work by Xuan et al. [41] proposed a joint image translation and reconstruction method which additionally accounts for misalignment between the reference and reconstructing images, with a follow-up work leveraging optimal transport theory [42]. Levac et al. [16] formulate guided reconstruction as a highly general Bayesian maximum a posteriori estimation problem and solve it iteratively via Langevin update steps, using an image-domain diffusion model as the score function of the prior distribution. While these are promising directions, we propose an alternative approach that decomposes the multi-contrast problem in a more intuitive way – first, into two sub-problems, namely image modeling and image reconstruction; and second, within the image-domain model, the multi-contrast information is decomposed into content and style. This two-level decomposition results in a highly modular reconstruction algorithm with a built-in explanatory framework where (a) the guidance mechanism is a simple content-replacement operation, (b) the discrepancy between the supplied reference content and the true content of the target image represents a meaningful error term which can be minimized, and (c) the optimal content-encoding capacity of the model for a given two-contrast image dataset indicates the amount of shared structure available to be learned in this data and used in the reconstruction task.
III Methods
III-A Reconstruction Problem
III-A1 Undersampled MRI reconstruction
Given a set of $P$ acquired k-space samples $y\in\mathbb{C}^{P}$ and the MRI forward operator $A\in\mathbb{C}^{P\times Q}$, CS reconstruction of the image $x\in\mathbb{C}^{Q}$ with $Q$ voxels is given as
$\min_{x}||Ax-y||_{2}^{2}+\lambda||\Psi x||_{1},$ (1)
where $\Psi$ is some sparsifying transform (e.g. wavelet) and $\lambda$ is the regularization strength. A commonly used algorithm to solve this optimization problem is ISTA, which iteratively applies the following two update steps:
$r^{k}\leftarrow\Psi^{H}\mathrm{soft}(\Psi x^{k-1};\lambda),$ (2)
$x^{k}\leftarrow r^{k}-\eta A^{H}(Ar^{k}-y).$ (3)
Eq. (2) performs soft-thresholding in the transform domain, thereby reducing the incoherent undersampling artifacts in image $x^{k-1}$, whereas (3) enforces soft data consistency on image $r^{k}$ by taking a single gradient descent step over the least-squares term, controlled by step size $\eta$.
III-A2 Plug-and-play denoiser
Plug-and-play methods replace the analytical operation of (2) with off-the-shelf denoisers. A CNN-based denoiser is of special interest as it incorporates a learning-based component into iterative reconstruction. Given a CNN model $f_{\theta}$ trained to remove i.i.d. Gaussian noise from an image, PnP-CNN [23] modifies Eq. (2) to
$r^{k}\leftarrow f_{\theta}(x^{k-1}).$ (4)
III-A3 Plug-and-play content consistency operator
In guided reconstruction, a spatially aligned reference $x^{\mathrm{ref}}_{1}$ is available, which captures the same underlying semantic content as the target reconstruction. Inspired by the PnP design, we cast the problem of incorporating prior information from $x^{\mathrm{ref}}_{1}$ into the reconstruction iterate $x_{2}^{k-1}$ as enforcing a hard consistency between this image iterate and its semantic content estimated from the reference. We propose a content consistency operator $g_{M}(\cdot;c)$ such that
$r_{2}^{k}\leftarrow g_{M}(x_{2}^{k-1};c),$ (5)
where a content/style model $M$ decomposes $x_{2}^{k-1}$ into content and style, followed by a replacement of this corrupted content with high-quality content $c$ derived from $x^{\mathrm{ref}}_{1}$ and composing the improved image $r_{2}^{k}$ from it. Before formally defining this operator and developing the reconstruction algorithm, we discuss the steps required to learn the content/style model $M$.
III-B Content/Style Modeling
Given an image dataset of two MR contrasts we make the general assumption that there exists an underlying contrast-independent structure which, influenced by arbitrary contrast-specific factors, is rendered as the contrast images.
III-B1 Unpaired pre-training of MUNIT
We formulate our content/style model based on the MUNIT framework [22]. We define an image domain as the set $\mathcal{X}_{i}$ of images of a certain contrast comprising the dataset, where $i\in\{1,2\}$. “Content” $c\in\mathcal{C}$ is defined here as the underlying contrast-independent structure and is represented as a set of feature maps, whereas “style” $s_{i}\in\mathcal{S}_{i}$ corresponds to the various modes of variability in one domain which cannot be explained by the other, e.g. global effects of acquisition settings, contrast-specific tissue features, etc., and is represented as a low-dimensional vector. MUNIT posits the existence of functions $G_{i}^{*}:\mathcal{C}\times\mathcal{S}_{i}\rightarrow\mathcal{X}_{i}$ and their inverses $E_{i}^{*}=(G_{i}^{*})^{-1}$, and learns them jointly via unpaired training, given samples from marginal distributions $p(x_{i})$. In practice, the encoder $E_{i}$ is split into content encoder $E_{i}^{c}$ and style encoder $E_{i}^{s}$. Thus, the content/style model is specified as $M=\{E_{1}^{c},E_{2}^{c},E_{1}^{s},E_{2}^{s},G_{1},G_{2}\}$.
The MUNIT loss function is comprised of 4 terms:
$\mathcal{L_{\mathrm{MUNIT}}}=\mathcal{L}_{\mathrm{GAN}}+\alpha_{1}\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}+\alpha_{2}\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}+\alpha_{3}\mathcal{L}_{\mathrm{style}}^{\mathrm{self}},$ (6)
where
$\begin{split}\mathcal{L}_{\mathrm{GAN}}=\>&\mathbb{E}_{x_{1}\sim p(x_{1}),s_{2}\sim q(s_{2})}[(1-D_{2}(G_{2}(E_{1}^{c}(x_{1}),s_{2})))^{2}]+\\ &\mathbb{E}_{x_{2}\sim p(x_{2}),s_{1}\sim q(s_{1})}[(1-D_{1}(G_{1}(E_{2}^{c}(x_{2}),s_{1})))^{2}],\end{split}$ (7)
$\begin{split}\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}=\>&\mathbb{E}_{x_{1}\sim p(x_{1})}[||x_{1}-G_{1}(E_{1}^{c}(x_{1}),E_{1}^{s}(x_{1}))||_{1}]+\\ &\mathbb{E}_{x_{2}\sim p(x_{2})}[||x_{2}-G_{2}(E_{2}^{c}(x_{2}),E_{2}^{s}(x_{2}))||_{1}],\end{split}$ (8)
$\begin{split}\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}=\>&\mathbb{E}_{c_{1}\sim p(c_{1}),s_{2}\sim q(s_{2})}[||c_{1}-E_{2}^{c}(G_{2}(c_{1},s_{2}))||_{1}]+\\ &\mathbb{E}_{c_{2}\sim p(c_{2}),s_{1}\sim q(s_{1})}[||c_{2}-E_{1}^{c}(G_{1}(c_{2},s_{1}))||_{1}],\end{split}$ (9)
$\begin{split}\mathcal{L}_{\mathrm{style}}^{\mathrm{self}}=\>&\mathbb{E}_{c_{1}\sim p(c_{1}),s_{2}\sim q(s_{2})}[||s_{2}-E_{2}^{s}(G_{2}(c_{1},s_{2}))||_{1}]+\\ &\mathbb{E}_{c_{2}\sim p(c_{2}),s_{1}\sim q(s_{1})}[||s_{1}-E_{1}^{s}(G_{1}(c_{2},s_{1}))||_{1}].\end{split}$ (10)
$\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}$ is the image recovery loss, which promotes preservation of the image information in the latent space. $\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}$ and $\mathcal{L}_{\mathrm{style}}^{\mathrm{self}}$ are the content and style recovery losses, respectively, which specialize parts of the latent code into content and style components. $\mathcal{L}_{\mathrm{GAN}}$ is the adversarial loss, which enables unpaired training by enforcing distribution-level similarity between synthetic and real images via discriminators $D_{1}$ and $D_{2}$. $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are hyperparameters.
III-B2 Paired fine-tuning
While the pre-training stage learns useful content/style representations, the model can be adapted to our task by improving its content preservation. To this end, we propose a paired fine-tuning (PFT) stage, leveraging a modest amount of paired data. Given samples from the joint distribution $p(x_{1},x_{2})$, our fine-tuning objective is given as
$\mathcal{L}_{\mathrm{PFT}}=\mathcal{L}_{\mathrm{GAN}}+\beta_{1}\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}+\beta_{2}\mathcal{L}_{\mathrm{image}}^{\mathrm{cross}}+\beta_{3}\mathcal{L}_{\mathrm{content}}^{\mathrm{cross}},$ (11)
where
$\begin{split}\mathcal{L}_{\mathrm{image}}^{\mathrm{cross}}=\>&\mathbb{E}_{\{x_{1},x_{2}\}\sim p(x_{1},x_{2})}[\\ &||x_{2}-G_{2}(E_{1}^{c}(x_{1}),E_{2}^{s}(x_{2}))||_{1}+\\ &||x_{1}-G_{1}(E_{2}^{c}(x_{2}),E_{1}^{s}(x_{1}))||_{1}],\end{split}$ (12)
$\begin{split}\mathcal{L}_{\mathrm{content}}^{\mathrm{cross}}=\mathbb{E}_{\{x_{1},x_{2}\}\sim p(x_{1},x_{2})}[||E_{1}^{c}(x_{1})-E_{2}^{c}(x_{2})||_{1}].\end{split}$ (13)
$\mathcal{L}_{\mathrm{image}}^{\mathrm{cross}}$ is a pixel-wise image translation loss, which provides image-level supervision, and $\mathcal{L}_{\mathrm{content}}^{\mathrm{cross}}$ is a paired content loss, which penalizes discrepancy between the contents. $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are an additional set of hyperparameters.
III-B3 Network architecture and content capacity
Following the original paper [22], our content encoders $E_{i}^{c}$ consist of an input convolutional layer potentially followed by strided downsampling convolutions, and finally a series of residual blocks. Style encoders $E_{i}^{s}$ consist of input and downsampling convolutions followed by adaptive average pooling and a fully-connected layer that outputs the latent vector. Decoders $G_{i}$ follow a similar structure as the content encoders except in reverse. Style is introduced into the decoder via AdaIN operations which modulate the activation maps derived from the content. We observe that in this architecture, the ratio between the content and image resolutions reflects the level of local structure one expects to be shared between the two domains. This (relative) content resolution is thus an inductive bias built into the model, which we refer to as the model’s content capacity. This concept is closely related to the general concept of the locality bias of image-to-image models [43]. A model with high content capacity has a large content resolution, allowing it to learn a rich content representation that strongly influences the output’s structure, which simultaneously restricts the style spaces to learn low-level global features.
Figure 2: A trained content/style model $M$ and the content consistency operator $g_{M}(\cdot;c)$ based on it. Images shown here are from the NYU brain DICOM dataset. Practical details are provided in Section IV. (a) A T1W/T2W image pair and the corresponding content maps. While these two contents generally agree, there is a notable discrepancy. (b) Synthetic T2W images generated from 16 style codes grid-sampled from $\mathcal{S}_{2}$, showing a smooth variation in image contrast and a roughly constant anatomical structure. (c) $g_{M}(\cdot;c)$ is applied on 3 cases where $x_{2}^{*}$ was corrupted with $R$=4 undersampling with different center fractions. Sampling more low-frequency lines leads to more accurate style estimate $\hat{s}_{2}$ and thus a better image $x_{2}^{\mathrm{cc}}$. Contours in $\mathcal{S}_{2}$ indicate the MAE of the synthetic images from that region. The red circle indicate a style estimation NMSE of 0.1.
III-C Iterative Reconstruction using Content Consistency
Given the content/style model $M$, we consider $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ as reference and target domains, respectively, in our guided reconstruction task, and define the content consistency operator $g_{M}(\cdot;c)$ from Eq. (5) as
$x^{\mathrm{cc}}_{2}=g_{M}(x^{\mathrm{us}}_{2};\hat{c}):=G_{2}(\hat{c},E^{s}_{2}(x^{\mathrm{us}}_{2})),$ (14)
where $x^{\mathrm{us}}_{2}$ is the image containing undersampling artifacts, $x^{\mathrm{cc}}_{2}$ is the content-consistent image, and $\hat{c}=E^{c}_{1}(x^{\mathrm{ref}}_{1})$ is the reference content. This operation improves $x^{\mathrm{us}}_{2}$ by simply replacing its aliased content with the content estimated from $x^{\mathrm{ref}}_{1}$, a rule which will later be softened with (16).
Note that this is a radical operation as it discards all structure contained in $x^{\mathrm{us}}_{2}$, retaining only a compact style code $\hat{s}_{2}=E_{2}^{s}(x^{\mathrm{us}}_{2})$. Let $x_{2}^{*}$ be the ground truth reconstruction with content $c^{*}$ and style $s_{2}^{*}$. The success of our content consistency operation depends on two conditions – (a) $\hat{c}$ is close to $c^{*}$ and (b) $\hat{s}_{2}$ is close to $s_{2}^{*}$. The first condition is roughly satisfied, as seen in Fig. 2a, because the model is explicitly trained to minimize content discrepancy, but more on this later. On the other hand, it is not obvious that the second condition should hold too and hence deserves a closer look. Assuming a high degree of shared local structure between the two domains, the optimal model has high content capacity (see Section III-B3). Thus, style would represent mostly low-level global image features, e.g. contrast variations, as indeed observe in Fig. 2b. It is a well known fact that image contrast is contained prominently in the center of the k-space. Hence, the estimate $\hat{s}_{2}$ can be made arbitrarily close to $s_{2}^{*}$ by sufficiently sampling the k-space center, as shown empirically in Fig. 2c.
Applying data consistency update (3) following $g_{M}(\cdot;\hat{c})$ in repetition yields an ISTA-based iterative scheme. Here, our content consistency update complements the data consistency update in the sense that while the latter forces the image estimate to be consistent with the given (measured) k-space data $y$, the former forces it to be consistent with the given (prior) content $\hat{c}$.
Algorithm 1 PnP-MUNIT
1:$y$, $A$, $x^{\mathrm{ref}}_{1}$, $M$, $\eta$, $\gamma$triangleright where $M=\{E_{1}^{c},E_{1}^{s},E_{2}^{c},E_{2}^{s},G_{1},G_{2}\}$
2:$k\leftarrow 0$
3:$x_{2}^{k}\leftarrow A^{H}y$triangleright Initialize Reconstruction
4:$c^{k}\leftarrow E_{1}^{c}(x^{\mathrm{ref}}_{1})$triangleright Initialize Content
5:repeat
6:     $k\leftarrow k+1$
7:     $r_{2}^{k}\leftarrow g_{M}(x_{2}^{k-1};c^{k-1})$triangleright Content consistency
8:     $x_{2}^{k}\leftarrow r_{2}^{k}-\eta A^{H}(Ar_{2}^{k}-y)$triangleright Data consistency
9:     $c^{k}\leftarrow c^{k-1}-\gamma\nabla_{c}||AG_{2}(c^{k-1},E_{2}^{s}(x_{2}^{k}))-y||_{2}^{2}$triangleright CR
10:until convergence
11:return $x_{2}^{k}$
The core assumption of our idealized content/style model is that contents $\hat{c}$ and $c^{*}$ are identical. However, in reality, this assumption will not hold, and a discrepancy between the two contents is to be expected. There are two possible sources of this discrepancy – (a) model-related, e.g. fundamental limits such as irreducible error [^1] If perfectly zero error was possible, one could perfectly predict the (distribution of) target image from the reference without any measurement. and practical issues like sub-optimal training, and (b) reference image-related, e.g. presence of artifacts independent of the target reconstruction. This content discrepancy, as observed empirically, is shown in Fig. 2a. During reconstruction, the error in $\hat{c}$ would limit the efficacy of the operator $g_{M}(\cdot;\hat{c})$, affecting reconstruction quality. While model-related discrepancy is partly tackled by PFT (Section III-B2), we now propose a content refinement (CR) process to correct for the remaining discrepancy in the reconstruction stage. Since we have no direct access to the true content $c^{*}$, but only to undersampled k-space measurements $y$, we aim at solving the following minimization problem in CR
$\min_{c}||AG_{2}(c,\hat{s}_{2})-y||_{2}^{2},$ (15)
starting from the initial point $\hat{c}=E^{c}_{1}(x^{\mathrm{ref}}_{1})$ and for a given style estimate $\hat{s}_{2}$. The augmented forward operator $AG_{2}(\cdot)$ is a composition of the linear MRI forward operator $A$ and the non-linear content/style decoder $G_{2}$, and it maps the content domain to the k-space domain. The error between the predicted k-space $AG_{2}(c,\hat{s}_{2})$ and the measured data $y$ serves as a proxy for content discrepancy, which can be computed and minimized during the reconstruction. We approximate the solution with a single gradient descent step
$c^{k}\leftarrow c^{k-1}-\gamma\nabla_{c}||AG_{2}(c^{k-1},\hat{s}_{2}^{k})-y||_{2}^{2},$ (16)
initialized as $c^{0}\leftarrow E^{c}_{1}(x^{\mathrm{ref}}_{1})$ and updated every $k^{\text{th}}$ iteration with step size $\gamma$ following the data consistency and content consistency updates. Hence, by aligning the content $c$ with k-space data $y$ and correcting the discrepancy, the CR module aligns content consistency updates with data consistency updates. With this additional component in place, we obtain our PnP-MUNIT reconstruction algorithm (Algorithm 1). Note that unlike content, the style of the image estimate need not be explicitly aligned with the k-space. The data consistency update applied on the image implicitly “corrects” its style and as the image converges, its style would converge as well.
IV Experimental Setup
IV-A Experiment Design
Without loss of generality, we considered the case of reconstructing T2W scans using T1W references and focused on head applications. We conducted 4 sets of experiments to comprehensively evaluate PnP-MUNIT – (a) simulations, (b) benchmark on NYU DICOM data, (c) benchmark on clinical multi-coil raw data, and (d) radiological evaluation.
TABLE I: Overview of the sequences in the LUMC datasets.
Sequence Params LUMC-TRA LUMC-COR
3D T1W TFE 2D T2W TSE 3D T1W TSE 2D T2W TSE
FA (deg) 8 90 80-90 90
TR (ms) 9.8-9.9 4000-5000 500-800 2000-3500
TE (ms) 4.6 80-100 6.5-16 90-100
ETL 200 14-18 10-13 17-19
Voxel size (mm) 0.98$\times$0.99 $\times$0.91 0.4$\times$0.54 0.59$\times$0.62 $\times$1.19 0.39$\times$0.47
Slice thick. (mm) 3 2
FOV (mm) 238$\times$191 $\times$218 238$\times$190 130$\times$238 $\times$38 130$\times$197
Num slices 50 15
TABLE II: LUMC data split. The splits were made at the subject-level, i.e. each subject belonged to exactly one subset. Splits containing only image-domain data are marked with I.
Dataset Split Subjects Sessions Scans (T1W / T2W)
LUMC-TRA model-trainI 295 418 360 / 415
model-valI 16 17 17 / 17
recon-val 18 21 21 / 21
recon-test 20 31 31 / 31
Total 339 487 429 / 484
LUMC-COR model-trainI 242 277 269 / 272
model-valI 18 18 18 / 18
recon-val 15 18 18 / 18
recon-test 16 17 17 / 17
Total 291 330 322 / 325
IV-A1 Simulation
The goal of our simulation experiments was to empirically test several properties of our algorithm under highly controlled settings. To this end, we used simulated T1W and T2W images based on BrainWeb phantoms [44]. BrainWeb provides 20 anatomical models of normal brain, each comprised of fuzzy segmentation maps of 12 tissue types. The 20 volumes were first split in 18:1:1 ratio for model training, validation, and reconstruction testing, respectively. T1W/T2W spin-echo scans were simulated using TE/TR values randomly sampled from realistic ranges. For testing the reconstruction, 2D single-coil T2W k-space data was simulated via Fourier transform and 1D Cartesian random sampling at various accelerations. We analyzed the effect of the content/style model’s disentanglement level and its content capacity on the reconstruction quality, the convergence of the PnP-MUNIT algorithm, and the effectiveness, robustness, and tuning of the CR module.
IV-A2 Benchmark on NYU DICOM dataset
In the NYU benchmark, we compared PnP-MUNIT against end-to-end reconstruction methods. The question we sought to answer was whether or not PnP-MUNIT, which requires only image-domain training data, can outperform methods that additionally require k-space training data. The NYU brain dataset [45, 46] includes raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR. Following [41], a paired DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans. All T1W scans were rigid-registered with the corresponding T2W scans. Single-coil T2W k-space data was simulated via Fourier transform with 1D Cartesian random sampling at $R\in\{2,3,4,5\}$ and added Gaussian noise of $\sigma=0.01\mathrm{max}(x_{2}^{*})$. The dataset was split into 4 subsets – model-train (200), model-val (27), recon-val (50), and recon-test (50). Images from the former two splits were used to pre-train, fine-tune, and validate the MUNIT model. On the other hand, end-to-end models were trained and validated on the aligned T1W, the simulated T2W k-space, and the T2W ground-truth from these two splits. The recon-val split was used to tune the PnP-MUNIT algorithm, and recon-test was the held-out test set. We benchmarked PnP-MUNIT against one well-known single-contrast network – MoDL [4] – and two recent multi-contrast networks – MTrans [32] and MC-VarNet [15].
IV-A3 Benchmark and ablation on LUMC multi-coil dataset
In our benchmark on clinical multi-coil data, our goal was to test PnP-MUNIT on a constrained real-world problem where only the image-domain data was available for training. This is often the case, as in clinical practice raw data is discarded after acquisition and only the final reconstructions are retained. Moreover, the T1W/T2W images were not fully subject-wise paired, representing a realistic case of data imbalance. Our in-house data consisted of brain scans of patients from LUMC, the use of which was approved for research purpose by the institutional review board. A total of 1669 brain scans were obtained from 817 clinical MR examinations of 630 patients acquired on 3T Philips Ingenia scanners. We focused on accelerating two T2W sequences – (a) 2D T2W TSE transversal and (b) 2D T2W TSE coronal. For guidance, two corresponding T1W sequences were used – (a) 3D T1W TFE transversal and (b) 3D T1W TSE coronal. The transversal and coronal protocols were considered as two separate datasets, namely LUMC-TRA and LUMC-COR. Table I shows an overview of the four sequences. Note that the in-plane resolution of the T1W scans was 1.3-2.5 times as low as that of the T2W scans, making these datasets more challenging for T1W-guided T2W reconstruction.
As with the NYU dataset, the LUMC datasets were split into model-train, model-val, recon-val, and recon-test. The former two splits contained only image-domain data, which was used to train and validate MUNIT. The MUNIT models were pre-trained on the full model-train splits, ignoring any pairing between T1W and T2W scans. Additionally, 20 subjects from the model-train split were designated for PFT where the pairing information was used in training and the reference scans were aligned via registration. The recon-val and recon-test splits additionally included multi-coil T2W raw data, which comprised 6-channel (LUMC-TRA) and 13-channel (LUMC-COR) k-space and coil sensitivity maps. This k-space was already undersampled (1D Cartesian random) at acquisition with clinical acceleration of $R$=1.8-2. We further undersampled it retrospectively to higher accelerations $R\in\{4,6,8,10\}$ by dropping subsets of the acquired lines. An overview of the data split is shown in Table II. Spatially aligned images required by all but the unpaired training set were obtained via rigid registration and the reference T1W images were resampled to the T2W resolution. During unpaired training, all images were resampled to the median T2W resolution.
Given the of absence of k-space training data, end-to-end reconstruction methods were not feasible as baselines. Hence, we compared PnP-MUNIT with only the feasible types of baselines, i.e. classical, semi-classical, and plug-and-play reconstruction and image-to-image translation. Among classical methods, we used the unguided L1-wavelet CS (CS-WT) and the guided STV-based CS (CS-STV) [7]. As an unguided plug-and-play baseline, we used PnP-CNN [23], and as a representative image translation baseline, we compared against MUNIT itself. Using MUNIT, deterministic image translation was approximated by combining the reference content with 200 randomly sampled T2W style codes and taking a pixel-wise mean of the resulting synthetic images to obtain a single synthetic image. As semi-classical guided baseline, we used (PROSIT) [39], which combines deterministic image translation with L2-regularized least-squares reconstruction. Additionally, as an ablation study for PnP-MUNIT, we ablated PFT and CR to assess their contribution, and finally, as an upper-bound for PnP-MUNIT representing zero content discrepancy, we disabled PFT and CR and used the ideal content $c^{*}$. In both NYU and LUMC benchmarks, we used 3 perceptual metrics for evaluation – SSIM, HaarPSI, and DISTS. While SSIM is used commonly, HaarPSI and DISTS are known to correlate better with visual judgment of image quality [47]. All three metrics are bounded in $[0,1]$ where $1$ represents perfect image quality. We conducted paired Wilcoxon signed-rank tests to measure statistical significance when comparing pairs of algorithms.
IV-A4 Radiological evaluation
Finally, as an extension to the LUMC benchmark, we conducted a radiological evaluation assessing the visual and diagnostic quality of PnP-MUNIT reconstructions. The study was conducted on a small sample of the LUMC-TRA recon-test set. With the help of a junior radiologist, we selected 3 cases with brain metastases. The evaluation comprised two parts, namely visual quality and pathology. Four visual quality criteria were used, namely sharpness, noise, artifacts, and contrast between gray and white matter and CSF. For pathology, we used three criteria, namely the number and sharpness of hyperintense areas within or surrounding metastases and the overall diagnostic quality of the scan for brain metastases. The images were scored using a five-point Likert scale [48] – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality. A diagnostic-quality reconstruction was defined as one that showed at least 90% of the metastases and scored at least “fairly diagnostic quality” on all other criteria. [^2] In clinical practice, while the remaining 10% of the metastases may not be detected in the reconstructed T2W scan, they would be prominently visible in the corresponding gadolinium-enhanced T1W scan, and hence would not go undetected. We evaluated PnP-MUNIT and PnP-CNN at 4 clinically realistic accelerations of $R\in\{3,4,5,6\}$. In total, 9 images per patient, including the (4$\times$2=8) reconstructions and the clinical ground truth (acquired at $R$=2 and reconstructed by the vendor software), were presented to a senior neuroradiologist who scored each image individually, blinded to the reconstruction method and $R$.
IV-B Implementation Details
We used the same general residual architecture for MUNIT encoders and decoders as the original paper, except with an additional layer at the end of the content encoder to produce content maps of given number of channels. We used 2 or 4 channels, depending on the content downsampling factor. The discriminators were implemented as multi-scale PatchGAN networks enabling them to locally assess the input images for realism at different scales. We used 1 scale in the BrainWeb simulations and 3 scales in the NYU and LUMC benchmark. We additionally conditioned the discriminators on foreground masks of the images to penalize background signal. To stabilize GAN training, we used spectral normalization [49]. For hyperparameters, we found that pre-training loss weight values $\alpha_{1}=\alpha_{2}=\alpha_{3}=1$ and paired fine-tuning values $\beta_{1}=\beta_{2}=\beta_{3}=1$ worked generally well for the in vivo datasets. In the benchmarks, the CR module parameter $\gamma$ of PnP-MUNIT was tuned per acceleration for each dataset and the number of iterations was set to 200.
In the simulations, all MUNIT models were pre-trained for 200k and were not fine-tuned for the sake of simplicity, unless stated otherwise in Section V-A. In both NYU and LUMC benchmarks, the MUNIT models were pre-trained for 400k iterations and fine-tuned for 50k iterations. All software was implemented in PyTorch, and image registration was performed using Elastix [50]. All training runs were performed on a compute node with an NVIDIA Quadro RTX 6000 GPU.
V Results
TABLE III: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices). PSNR worse than L1-wavelet CS is marked as .
Disentanglement Strength R=2 R=4
$\alpha_{2}=\alpha_{3}=10$ 15.14 $\pm$ 0.06 12.84 $\pm$ 0.05
$\alpha_{2}=\alpha_{3}=1$ 28.43 $\pm$ 0.07 26.13 $\pm$ 0.07
$\alpha_{2}=\alpha_{3}=0.1$ 29.84 $\pm$ 0.05 27.33 $\pm$ 0.05
$\alpha_{2}=\alpha_{3}=0.01$ 25.34 $\pm$ 0.06 23.03 $\pm$ 0.05
$\alpha_{2}=\alpha_{3}=0.001$ 18.17 $\pm$ 0.07 15.88 $\pm$ 0.05
TABLE IV: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices) at $R$=4. Optimal content capacity is shown in bold. PSNR worse than L1-wavelet CS is marked as .
Model Config Data Config
RefRes-1 RefRes-2 RefRes-4 RefRes-8
ContentRes-1 27.90$\pm$0.08 21.17$\pm$0.05 19.11$\pm$0.07 16.43$\pm$0.07
ContentRes-2 26.19$\pm$0.08 22.80$\pm$0.08 20.05$\pm$0.08 16.57$\pm$0.08
ContentRes-4 23.98$\pm$0.08 20.88$\pm$0.10 19.44$\pm$0.09 16.88$\pm$0.06
Figure 3: (a) Convergence of PnP-MUNIT and the baseline versions at $R$=4. (b) Evolution of the reconstruction shown with its error map.
Figure 4: Effectiveness of the CR module in resolving contrast-specific structure, which, in this case, was a simulated lesion.
V-A Experiments on Simulated MR Datasets
V-A1 Perturbation of disentanglement strength
PnP-MUNIT relies heavily on a sufficient disentanglement of content and style from the images. Here, we evaluated the effect of perturbing the disentanglement loss weights $\alpha_{2}$ and $\alpha_{3}$ (Eq. (6)) of MUNIT on the PnP-MUNIT reconstruction quality. Table III shows the reconstruction PSNR for different weight values. We observed that optimal disentanglement, and thus the best reconstruction, was achieved at $\alpha_{2}=\alpha_{3}=0.1$. At lower values, poor disentanglement of content and style in the model resulted in poor PnP-MUNIT reconstruction quality. This was expected since PnP-MUNIT relies on the assumption that the content representation is (sufficiently) contrast-independent, which was less enforced at these levels. At higher values, the training over-emphasized disentanglement while under-emphasizing the GAN and image recovery loss terms, thereby leading to worse preservation of image information in the latent space and thus, to worse reconstructions.
V-A2 Content resolution analysis
PnP-MUNIT relies on a reference image of sufficiently high resolution to provide guidance to the reconstruction. The goal of this experiment was to determine the effect of lowering the reference image resolution on the reconstruction quality and using the content/style model’s content capacity (i.e. the spatial resolution of the content relative to the images) to explain this effect. We simulated 4 datasets, denoted as RefRes-$n$, where $n\in\{1,2,4,8\}$ represents the reference domain downsampling factor. In the $n$=1 case, T1W/T2W images had the same resolution (as that of the underlying tissue maps), whereas in the subsequent cases, T1W images were blurred to contain only the lower $1/n$ frequency components, while maintaining the same spatial resolution. For each dataset, we trained 3 content/style models with content capacity denoted as ContentRes-$m$, where $m\in\{1,2,4\}$ is the content downsampling factor, which depends on the number of downsampling blocks in the networks. E.g. ContentRes-2 model produced content maps half the spatial resolution of the ContentRes-1 model.
Figure 5: The effect of misaligned reference on the reconstruction and the robustness of the CR module to it.
Figure 6: Reconstruction PSNR as a function of acceleration $R$ and k-space noise level $\sigma$ for different values of the $\gamma$ parameter. $\gamma$ controls the consistency of the content with the measured k-space.
Figure 7: Evaluation plots over recon-test subset of NYU DICOM dataset consisting of 550 slices.
Figure 8: Two examples from the NYU benchmark. (a) Comparison of all the methods at the highest acceleration of $R$=5. (b) Comparison between PnP-MUNIT and MC-VarNet at $R$=2 and $R$=4 on a pathological case where many features of the edema are T2W contrast-specific.
Table IV compares PnP-MUNIT reconstruction quality across these configurations. PFT and CR were disabled here for simplicity. We observe two trends. First, the reconstruction quality generally decreases from left to right, eventually dropping below L1-wavelet CS reconstruction, suggesting a decrease in the amount of shared local information contained in the reference contrast. Second, the optimal content capacity of the model decreases with the reference resolution in accordance with the actual amount of this shared information. In other words, this optimal content capacity corresponds to the amount of contrast-independent structure the model discovers in the dataset, e.g. RefRes-1 dataset had 4 times as much shared structure (due to twice as large optimal content resolution) as RefRes-2 and RefRes-4 datasets and hence, the RefRes-1 dataset was significantly more effective in guided reconstruction. On the other hand, Refsize-2 and RefRes-4 datasets had similar levels of shared structure despite the lower reference resolution in the latter, suggesting a more complex relationship between the reference resolution and the shared content. In the RefRes-8 case, PnP-MUNIT dropped to a similar level of quality as conventional CS, suggesting a lowerbound on the amount of shared information for PnP-MUNIT to be effective, specifically 1/42=1/16 times the local information of the full-resolution case.
V-A3 Convergence
We explored the convergence of PnP-MUNIT by comparing it with two variants. The first used true content $c^{*}$ of the ground truth T2W image, thus assuming zero content discrepancy and representing an upper-bound of PnP-MUNIT. The second used reference content $\hat{c}$ with CR step disabled, representing a lower-bound of PnP-MUNIT where the non-zero content discrepancy is not corrected. Here and in the following simulation experiments, we used the RefRes-1 dataset and the ContentRes-1 model additionally fine-tuned on 2 training volumes for 50k iterations. Fig. 3 shows the convergence curves and intermediate reconstructions. The upper-bound version, given $c^{*}$, converged in a single iteration, while with $\hat{c}$, the ablated version converged equally fast but to a sub-optimal solution. Enabling the CR step closed the gap with the upper-bound, although costing convergence rate.
V-A4 Effectiveness in resolving contrast-specific structure
In order to test the effectiveness of the CR module in resolving structure present in the target contrast but absent in the reference contrast, we simulated a lesion in the T2W image and performed reconstruction with and without the CR module. As shown in Fig. 4, the CR module was able to recover the lesion fully at $R$=2 and substantially at $R$=4.
V-A5 Sensitivity to misalignment of the reference
While PnP-MUNIT expects an aligned reference image at reconstruction time, the CR module should, in principle, correct for small misalignments. To test this, we simulated a 2 rotation in the reference image and performed reconstruction with and without the CR module, comparing also with PnP-MUNIT that used an aligned reference image. As shown in Fig. 5, although the reconstruction was sensitive to the misalignment in the absence of CR, it significantly improved with CR enabled.
Figure 9: Evaluation plots over recon-test subset of LUMC-TRA and LUMC-COR datasets consisting of 1366 and 200 slices, respectively.
Figure 10: Sample slice from the LUMC-COR recon-test set comparing PnP-MUNIT with PnP-CNN and MUNIT image translation.
V-A6 Tuning the CR step size
The hyperparameter $\gamma$ in Algorithm 1 controls the CR strength. Fig. 6 shows reconstruction quality as a function of acceleration and k-space noise level, where we observe that the optimal $\gamma$ decreased with the amount and quality of k-space data. This makes sense as enforcing agreement with the k-space gradually becomes less beneficial, thereby contributing less to reconstruction quality.
V-B Benchmark on NYU DICOM Data
Fig. 7 shows evaluation metrics for PnP-MUNIT and the end-to-end baselines on the NYU DICOM dataset. Despite being trained only on image-domain data, PnP-MUNIT largely outperformed MoDL ($p<0.05$ throughout, except SSIM at $R$=2 and DISTS at $R\in${2,3,4}) and MTrans ($p<$0.05), and was roughly comparable to MC-VarNet. Moreover, as shown in Fig. 8a, PnP-MUNIT could resolve certain fine details even at the high acceleration of $R$=5, which the baselines failed at. In the pathological case shown in Fig. 8b, PnP-MUNIT produced sharper reconstructions preserving the structure of the edema better than MC-VarNet despite lower metrics when many features of edema were absent in the reference image.
V-C Benchmark and Ablation on LUMC Multi-coil Data
The LUMC datasets represent a more challenging problem with real-world data constraints. Additionally, in the light of the content resolution analysis (Section V-A2), we empirically found the ContentRes-2 model configuration as optimal for the LUMC datasets, compared to NYU DICOM data for which ContentRes-1 was the optimal model configuration. This reflects the lower effective content in LUMC image data and hence a greater difficulty for guided reconstruction.
Fig. 9 plots the benchmark metrics for the LUMC test sets, which can be summarized in the following three trends. First, pure image translation with MUNIT was worse compared to single-contrast reconstruction, especially at lower acceleration factors (comparing with PnP-CNN, $p$$<$0.05 for all metrics and both datasets). Combining it with L2-regularized reconstruction in PROSIT improved SSIM and HaarPSI ($p$$<$0.05 for both datasets and accelerations), but not necessarily DISTS, suggesting that the the available complementary information was not fully utilized. PnP-MUNIT was consistently better than both MUNIT image translation and PROSIT ($p$$<$0.05 throughout for both cases). It also outperformed the conventional guided CS-STV ($p$$<$0.05 throughout, except SSIM at $R$=4 in LUMC-COR). Hence, using both measured k-space and reference scan via PnP-MUNIT was more beneficial than (a) using either one of them and (b) combining both using hand-crafted priors, suggesting that our approach maximally exploits the complementary information. Second, both PnP-CNN and PnP-MUNIT performed similarly at lower acceleration, except in terms of SSIM and HaarPSI at $R$=4 and $R$=6 on LUMC-COR where PnP-CNN was slightly better. At higher acceleration, PnP-MUNIT outperformed PnP-CNN ($p$$<$0.05 for all metrics and both datasets except SSIM at $R$=8 in LUMC-COR). Compared to PnP-CNN, PnP-MUNIT allowed up to 32.6% more acceleration for a given SSIM. [^3] Based on linear interpolation of LUMC-TRA SSIM, PnP-MUNIT and PnP-CNN allowed $R$=10 and $R$=6.7, the maximum difference in $R$, at median SSIM=0.906. Hence, a difference of 32.6%. Third, introducing PFT and CR into the ablated PnP-MUNIT improved the reconstructions ($p$$<$0.05 throughout in both cases except in DISTS for the latter at $R$=10 in LUMC-TRA). The contribution of PFT was almost constant across $R$, whereas that of CR decreased with $R$, which was expected since CR depends on the measured k-space data to refine the content.
Fig. 10 shows an LUMC-COR example representative of the first two trends. MUNIT image translation produced severe anatomical defects and predicted false structure in the lower region of the image where the ground truth, in fact, contains low signal. At $R$=4, PnP-MUNIT was visually similar to PnP-CNN, with the main difference being a blur effect in PnP-CNN and mild texture artifacts in PnP-MUNIT. At $R$=10, PnP-CNN reconstruction was non-viable, containing severe blur and artifacts (green arrows). On the other hand, PnP-MUNIT did not degrade much from the $R$=4 case and most fine structures were sharply resolved (green arrows). However, an interesting failure mode of PnP-MUNIT at high acceleration was the subtle localized distortions in the anatomy (pink arrows).
Figure 11: Averaged results of the radiological evaluation.
V-D Radiological Evaluation
Figure 11 plots the radiological evaluation result. At $R$=3, both algorithms produced at least “fairly diagnostic” reconstructions and allowed the detection of more than 90% of metastases, a difference of 33.3% of k-space samples over the clinical reconstructions of $R$=2. PnP-CNN was better than PnP-MUNIT at $R\in\{3,4\}$ in terms of sharpness, contrast, and the pathology criteria. At $R$=5, PnP-MUNIT matched or exceeded PnP-CNN in terms of visual quality, although being worse in terms of pathology criteria. At $R$=6, PnP-CNN sharply dropped to non-diagnostic level, whereas PnP-MUNIT was better, especially in terms of sharpness and contrast as well as in overall diagnostic quality.
VI Discussion
In this work, we modeled the two-contrast MR image data in terms of two latent generative factors – content representing the contrast-independent structure and style representing the contrast-specific variations. Our data-driven definitions of content and style, though seemingly related to the MR physics-based representations of quantitative maps and acquisition-related factors, are rather nebulous compared to precisely defined physical concepts. For instance, MR quantitative maps are theoretically objective representations of tissues in terms of physical variables such as relaxation times and are ideally independent of the contrast-generating sequence parameters. On the other hand, our content and style representations are only defined statistically in terms of the given two contrasts and are, thus, neither as objective nor as independent in the ideal physical sense. That being said, we would argue that there is potential value in augmenting the data-driven content/style model with MR physical models, e.g. by constraining the learned content to represent physically meaningful anatomical properties, introducing elements of a physical model (e.g. Bloch equation) into the decoder network, etc., to enhance interpretability and reconstruction quality.
Optimal content capacity was defined as a quantity representing the amount of shared local information contained in a two-contrast dataset. In terms of this quantity, we analyzed the effect of reference image resolution on reconstruction quality, obtaining a lowerbound for PnP-MUNIT to be effective (Section V-A2). This quantity would depend on more fundamental factors such as MR sequence types, which we did not investigate here. While we limited our experiments to T1W and T2W sequences, PnP-MUNIT is applicable to any pair of contrasts, with the image quality limited mainly by the amount of structural information shared between them. Future work could help empirically determine which sequence pairs are more amenable to guided reconstruction than others. Note that for a different pair of contrasts, say PD-weighted and T2W, a fresh content/style model needs to be trained since content and style are defined in the context of the specific contrast pair. It is, in principle, also possible to model the content and style for more than two contrasts. Given number of contrasts $N$, content is defined as the local structure underlying all $N$ contrasts, while style encodes intra-contrast variations of each individual contrast. This model would include $N$ sets of encoders and decoders and would require an (unpaired) image-domain training dataset of these $N$ contrasts. Subsequently, our PnP-MUNIT framework can be extended to use multiple reference contrasts to guide a given target contrast. The specific question to investigate then would be – given $P$ reference contrasts (where $1<P<N$), how to aggregate their contents such that the prior structural information about the target image is maximized?
Content discrepancy was characterized as a quantity representing the gap between the true content of the target image and our estimation of this content, which is another limiting factor in PnP-MUNIT reconstruction. In MUNIT, the decoders model the (forward) generative process that maps the latent variables (i.e. content and style) to the observable variable (i.e. images). Ideally, the encoders must be a perfect inverse of this forward process. However, estimating the multi-channel contrast-independent content from a single image (using the content encoders) is a challenging (and perhaps ill-conditioned) inverse problem. Hence, some errors are to be expected in the content estimated from the reference image. Moreover, the reference image may contain unfavorable differences from the target image, such as misalignment or artifacts, which introduce more errors. Therefore, content discrepancy is expected to be a non-zero quantity. We proposed two complementary ways to minimize its effect on the reconstruction. While paired fine-tuning improves the model using a small amount of paired image-domain training data, content refinement (CR) aims to correct the remaining discrepancy during reconstruction using the measured k-space. A drawback of the CR module, however, is slower convergence (Section V-A3). For example, the run-time for PnP-MUNIT on the GPU was around 17 seconds for a slice of matrix size 349$\times$284, as compared to 6 seconds for CS-WT, 1.4 seconds for PnP-CNN, and less than 0.1 second for MC-VarNet. This latency issue could be mitigated by approximating the iterative process with an unrolled network design, although perhaps at a cost of the generalizability offered by the plug-and-play design. The resulting trade-off between model latency and generalizability would be interesting to explore in the future.
In terms of training data requirements, PnP-MUNIT can make use of the larger amounts of the image data available and can be applied to situations where end-to-end methods are infeasible since it does not rely on k-space training data. Moreover, unpaired image-domain pre-training of the content/style model boosts the practical applicability of PnP-MUNIT, e.g. in the case of LUMC-TRA where the T1W/T2W data imbalance was considerable. As indicated in our NYU DICOM benchmark, PnP-MUNIT performs similarly or better than state-of-the-art end-to-end methods. In our LUMC multi-coil benchmark, we limited our baselines to those algorithms that were feasible given the data constraint. As future work, a more comprehensive study should be conducted with a broader range of methods and evaluation criteria, e.g. comparing generalizability across accelerations and sampling patterns with multi-contrast unrolled networks, given a fixed budget of paired training data. Contrary to the training stage, we strictly assumed the spatial alignment of the reference and target images in the reconstruction stage. The CR update, which can implicitly correct minor registration errors, would likely break down at the typically observed levels of patient motion. A potential solution is to incorporate an online registration step to explicitly and efficiently correct for arbitrary inter-scan motion, thereby further improving practical applicability.
In terms of reconstruction quality, we observed in Section V-C that the true added value of PnP-MUNIT was at the high acceleration factors ($R$=8 and $R$=10) where the k-space data is scarce and the reference information becomes more valuable. However, the risk of model hallucinations increases at these accelerations, raising concerns about the accuracy of the anatomy represented in the image. Given the advantages of content/style decomposition, it may be possible to leverage the contrast-independence property of the content to automatically detect local hallucinations in the reconstruction and to define a corrective process to minimize it. This is another topic for future work. At lower accelerations, on the other hand, we observed that PnP-MUNIT often produced slightly lower metrics compared to PnP-CNN, e.g. in LUMC-COR. This was also observed in the radiological evaluation on LUMC-TRA samples. This is counter-intuitive at first glance since PnP-MUNIT has access to additional side-information and should, in principle, perform at least as good as PnP-CNN. However, this can be explained by the fact that at lower accelerations, the content/style model becomes the limiting factor for reconstruction quality. This is because unlike the CNN denoiser (Eq. (4)) of PnP-CNN, the content consistency operator (Eq. (14)) of PnP-MUNIT radically changes the image and hence, this operation is sensitive to the model’s overall performance. The model performance is in turn influenced by the training dataset and the network architecture. First, regarding data, the T1W scans in LUMC datasets had an in-plane resolution 1.3-2.5 times lower than that of the T2W scans, limiting the quality of the model (as indicated by its low optimal content capacity) and the overall value of guidance. In the long run, this may be solved by using 3D reference and target sequences of matched-resolution to optimize guidance. Second, regarding model design, correcting for content discrepancy using the CR module relies on computing gradients through the decoder $G_{2}$ (Eq. (16)). If $G_{2}$ is highly non-linear and the reference content is not sufficiently close to the optimum, CR may converge to a sub-optimal point leading to texture artifacts in the reconstruction, as seen in Fig. 10 at $R$=4. In our proof-of-concept, we used the same general model architecture proposed in [22]. Better designs that are more conducive to CR may exist. Hence, future efforts must considerably focus on improving the content/style model using specialized architecture and regularization.
VII Conclusion
In this work, we introduced PnP-MUNIT, a modular approach to multi-contrast reconstruction combining content/style modeling with iterative reconstruction, which offers the reconstruction quality of end-to-end methods under stronger training data constraints. At its core is the content consistency operation, which provides regularization at the level of the image’s semantic content. We defined two quantities that determine the efficacy of this operation, namely optimal content capacity and content discrepancy, and provided several ways of maximizing this efficacy. On real-world clinical data, PnP-MUNIT provided up to 32.6% more acceleration over PnP-CNN for given SSIM, enabling sharper reconstructions at high accelerations. In the radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-MUNIT produced diagnostic-quality images at $R$=3, enabling 33.3% more acceleration over clinical reconstructions. To progress towards practical implementation of our proof-of-concept, future work will focus mainly on improving the content/style model and reconstruction latency, tackling the problem of model hallucinations, and incorporating online registration into the reconstruction.